she biked for 48, 49, and 65 miles on the first three days. How many miles does she need to bike on the last day so that her average is 54 miles per day
Found 2 solutions by rfer, macston:
Answer by rfer(16322)

(48+49+65+x)/4=54
(162+x)/4=54
162+x=54*4
x=216-162
x=54

Answer by macston(5194)

.
To average 54 miles over 4 days, her total mileage must be:
4(54 miles)=216 miles
After 3 days her total is:
48mi+49mi+65mi=162 miles
On the fourth day, her mileage must be:
216mi-162mi=54 miles